if 8,x,y,27 are in GP. Find x and y?
Answer by richwmiller(17219)

8*a=x,x*a=y,y*a=27
8*a^3=27
a = 3/2, x = 12, y = 18
8*3/2=12
12*3/2=18
18*3/2=27
BTW There are at least two complex solutions too.